y is inversely proportional to x. y=30 when x=5.
(a) Find the constant of proportionality.
(b) Determine the value of y when x=15.

Answer by stanbon(75887)

y is inversely proportional to x. y=30 when x=5.
y = k/x
30 = k/5
k = 150
-------------------
(a) Find the constant of proportionality.
k = 150
--------------------
(b) Determine the value of y when x=15.
y = 150/15
y = 10
--------------
Cheers,
Stan H.
----------------